A plane drops a package to a party of explorer. If the plane is travelling horizontally at 40 meter per second and displacement is 100m above the ground. Where does the package strike the ground relative to the point at which itis relative to the point at which it is released.
1
Expert's answer
2021-08-29T16:56:44-0400
"t=\sqrt{\frac{2\cdot100}{9.8}}=4.518 s\\x=40(4.518)=180.7\ m"